OpenStudy (lgbasallote):

another question from my algebra book ;) an openstudy user gets his first fan 4 months after he joined. he gets to 100 fans 2 months after this. then he gets to 200 fans 19 days after. He gets to 300 fans 15 days after. Then he gets to 400 fans 29 days after. He reaches 500 fans 20 days after. create a function/equation representing his progress

Parth (parthkohli):

lol

Parth (parthkohli):

that's you ;p

OpenStudy (lgbasallote):

lol you're fast

OpenStudy (lgbasallote):

you came here before i did

Parth (parthkohli):

There's no function I see here

OpenStudy (lgbasallote):

then how can i represent this?

OpenStudy (lgbasallote):

and no this aint me..im not 500 fans :p

Parth (parthkohli):

Though, f(1) = 120 f(100) = 180 f(200) = 199 f(300) = 214 f(400) = 243 f(500) = 263 You can create an array function

OpenStudy (lgbasallote):

i see

Parth (parthkohli):

f(x) = {120 if x = 1 {180 if x = 100 etc.

Parth (parthkohli):

There's no common thing I see between them lol

Parth (parthkohli):

And good job, you've kept the records xD

OpenStudy (lgbasallote):

this is an even-number in my book so it doesnt give answers :/

Parth (parthkohli):

So is it really about OpenStudy or you modified it a bit?

OpenStudy (lgbasallote):

openstudy..i wrote it remember :p

OpenStudy (lgbasallote):

so summary...a progress has to be consistent for it to be function-ed?

Parth (parthkohli):

How does your book know about OS? o.o

Parth (parthkohli):

Yeah, it has to have something in common.

OpenStudy (lgbasallote):

i wrote it :p

Parth (parthkohli):

This, my dear sir, is completely random :P

OpenStudy (lgbasallote):

*my* book

OpenStudy (lgbasallote):

making any sense?

Parth (parthkohli):

oh lol

Parth (parthkohli):

YOUR book xP

OpenStudy (lgbasallote):

lol

Parth (parthkohli):

functions have infinite number of notations, but this array thing is the one I'm telling you

Parth (parthkohli):

Yours might be different tho

Parth (parthkohli):

How do you make a { in latex?

OpenStudy (lgbasallote):

dunno

Parth (parthkohli):

Oh yeah, Sat knows. Let me go to the \(\LaTeX\) group.

OpenStudy (lgbasallote):

\[\huge \text{{}}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large\left\{ \begin{matrix}f(x) = x^2 - 1 \\ f(x) = x^2 + 1 \end{matrix}\right.}\)

Parth (parthkohli):

Yeah it was easy

OpenStudy (lgbasallote):

dont spam my post lol